# Are LLMs Prescient? A Continuous Evaluation using Daily News as the Oracle

### v1

@@ 4.1 Experimental Setup
In addition to a closed-book evaluation, we explore the constrained open-book setting: how access to news articles up to different time cutoffs influences LLM performance using RAG (Lewis et al.,, 2020). We introduce the concept of the RAG cutoff (R-Cutoff), which limits the latest accessible date for retrieving articles. To prevent the models from leveraging information beyond the resolution date, for any question with a resolution date ($d_{\text{res}}$), the accessible articles span from January 1st, 2019 (the start of our news corpus) up to whichever comes first between the day before the resolution date and the RAG cutoff date ($d_{\text{R-Cutoff}}$). Formally, the accessible date range is $[01/01/2019,\min(d_{\text{res}}-1,d_{\text{R-Cutoff}}))$. Following prior work (Jin et al.,, 2021; Zou et al.,, 2022; Zhang et al.,, 2024), we employ BM25 (Robertson et al.,, 1995) as the retriever and select the top 5 articles relevant to each question. We truncate each retrieved article to a maximum length of 512 words. These articles are then incorporated into the input prompts to serve as additional information.
@@
We further include a setting where models are provided direct access to the gold article, from which the question is generated. This transforms the forecasting questions into reading comprehension ones, which can also access LLMs’ general question-answering capabilities. Achieving high accuracy here ensures that the questions from our Daily Oracle dataset are answerable.
@@
Accuracy score is used as the evaluation metric. Though LLMs are tested daily, to show clearer trends, we plot the monthly performance in Figure 4, and apply a 5-month moving average to smooth the curve. We also report yearly averages and average year-over-year (YoY) accuracy change before and after models’ knowledge cutoff dates in Table 3. Additionally, despite prompting the models to avoid responses like “I cannot predict the future” and instead provide definitive answers, there are cases where such refusals still occur. The rejection rates are provided in the Appendix B.2, and these cases are counted as incorrect to ensure comparability across model results.
@@ 4.2 Main Results
Figure 4 and Table 3 present our primary results for the closed-book setting. The “Avg” column in Table 3 shows the average YoY accuracy change of all months, revealing a clear degradation in performance over time across all models on both TF and MC questions. When comparing accuracies from the beginning to the end of the evaluation period, we observe that, on average, the models’ performance declines by 20.14% on TF questions (from 64.68% to 51.65%) and by 23.26% on MC questions (from 58.30% to 44.74%). This indicates that while LLMs demonstrate certain abilities to understand real-world events and make predictions, they struggle to maintain these abilities.
Figure 4: Results for the closed-book setting. We plot the 5-month moving average accuracy for True/False (TF) and Multiple Choice (MC) questions across different LLMs.
Notably, the average YoY accuracy declines provide further insight. Before the knowledge cutoff, the average YoY decline across all models was relatively moderate. However, post-knowledge cutoff, we observe steeper declines in many models, with GPT-4 showing the most drastic drop in MC performance, declining by 18.47%, compared to just 4.23% before the cutoff. This contrast highlights that while LLMs manage to retain a baseline of past knowledge with small degradation, their ability to forecast future events deteriorates much more rapidly as they move beyond their training data, struggling with temporal generalization.
Among different models, Claude-3.5-Sonnet (Anthropic,, 2024) significantly outperforms all others, while GPT-4 excels in MC questions but its performance in TF is not as remarkable as in MC. GPT-3.5, Qwen-2-7B (Yang et al.,, 2024) and Llama-3-8B (Dubey et al.,, 2024) show smaller temporal declines than GPT-4 in both TF and MC questions. Interestingly, Mistral-7B (Jiang et al.,, 2023) and Mixtral-8x7B (Jiang et al.,, 2024) show the most pronounced drops in TF accuracy, with scores falling below the random baseline 50% due to increased answer refusals, as shown in the Appendix B.2. Gemma-2-2B (Team et al.,, 2024) exhibits the most consistent performance with the smallest average YoY decline, likely due to its more recent knowledge cutoff date.
Table 3: For various LLMs with different knowledge cutoffs (K-Cutoffs), we show the yearly average accuracy (calculated as the average across months) from 2020 to 2024, along with the average year-over-year (YoY) accuracy change (%) before the knowledge cutoff date (Pre-Cutoff), after the knowledge cutoff date (Post-Cutoff), and the overall average YoY accuracy change across all months (Avg) on the Daily Oracle dataset.
2020 2021 2022 2023 2024 Pre-Cutoff Post-Cutoff Avg
TF Claude-3.5-Sonnet Apr 2024 81.21 79.88 78.05 74.38 66.25 -4.77 -11.97 -4.47
GPT-4 Apr 2023 69.68 66.41 60.36 60.54 57.88 -5.83 -1.96 -4.21
GPT-3.5 Sept 2021 62.86 60.12 59.36 57.11 57.80 -4.33 -3.43 -2.11
Mixtral-8x7B Unknown 57.83 52.69 43.09 39.34 36.29 -10.93
Mistral-7B Unknown 57.57 54.65 48.22 41.35 41.89 -7.67
Llama-3-8B Mar 2023 65.06 64.24 62.35 58.68 56.44 -1.95 -6.5 -3.23
Qwen-2-7B Unknown 62.39 60.15 57.67 53.39 53.14 -3.86
Gemma-2-2B Jul 2024 58.71 59.31 57.64 56.61 55.87 -1.41 -5.28 -0.97
MC Claude-3.5-Sonnet Apr 2024 76.86 77.67 74.32 69.37 61.79 -6.26 -12.82 -4.83
GPT-4 Apr 2023 70.59 70.62 66.75 56.40 50.96 -4.23 -18.47 -7.48
GPT-3.5 Sept 2021 50.27 50.36 44.43 41.43 42.32 0.14 -0.46 -4.25
Mixtral-8x7B Unknown 57.38 56.97 50.76 47.10 45.09 -5.37
Mistral-7B Unknown 50.07 52.36 48.06 44.40 44.08 -2.56
Llama-3-8B Mar 2023 52.44 54.18 50.66 47.94 45.97 -2.21 -1.25 -3.01
Qwen-2-7B Unknown 55.28 55.93 53.44 49.77 47.94 -3.04
Gemma-2-2B Jul 2024 47.87 50.71 46.81 45.20 43.65 -4.46 -2.33 -1.73
@@
In Figure 5, we present the results of the constrained open-book setting, with Mixtral-8x7B on TF questions and Llama-3-8B on MC questions across different RAG cutoff dates. [^3] Refer to Appendix B.4 for results of other models in the constrained open-book setting. For Mixtral-8x7B, as the RAG cutoff dates extend to closer to the resolution dates, we observe a clear improvement in performance, indicating the model benefits from increasingly updated information retrieval. However, there are noticeable performance drops immediately after each RAG cutoff date when compared to providing information up to the day before the resolution date (with the exception of the March 2024 cutoff). This highlights the importance of keeping up-to-date information for optimal RAG performance. Interestingly, RAG does not uniformly enhance performance. Llama-3-8B may perform worse than the closed-book setting when the RAG cutoff is prior to the knowledge cutoff dates, suggesting outdated information may negatively impact performance. Conversely, for more recent RAG cutoff dates that extend beyond the knowledge cutoff, significant performance improvements are observed (as illustrated by the curves with cutoffs in September 2023 and March 2024). Notably, across all different RAG cutoffs, the overall performance decline pattern persists, likely due to outdated internal representations and the model’s inherent knowledge limitations.
@@
Figure 6 shows that when given access to the gold articles from which the questions are generated, LLM performance can be improved significantly to around 90%, demonstrating the answerability of Daily Oracle. [^4] Results for GPT-3.5 are provided and discussed in Appendix B.3, as this older model performs relatively poorly and including it on the same scale would obscure the trends of other models. However, most of the models still show declining trends. This is noteworthy because, ideally, LLMs are expected to achieve consistent accuracy regardless of the article’s publication date when answers are directly accessible. However, the outdated representations hinder their ability to consistently generate correct answers, even in a reading comprehension setting.
@@ 4.3 Discussion
LLMs’ Performance Evolution Across Time.
@@
We observe several LLMs’ performance evolution patterns in Figure 4: (1) Gradual Decline in the Recent Past: In the months before the knowledge cutoff date, which we call the recent past, we observe a gradual decline in model performance, as seen in Llama-3-8B, GPT-4, and Claude-3.5-Sonnet, likely due to a lack of representation of recent news in the training data. (2) Rapid Decline in the Near Future: In the near future, which we define as the months following a model’s knowledge cutoff date, sharp performance drops are observed in several models in MC questions. For instance, the decline in Claude-3.5-Sonnet and GPT-4 accelerates soon after their knowledge cutoffs. Most of the models, however, do not lose all the predictive power at once, as evidenced by the further decline into the farther future.
@@
We explore this further by analyzing the slope of accuracy as a function of time. In Figure 7, we show how the slope changes as we fit a regression to an increasingly larger window of data, until we reach the full set of accuracies. Specifically, using the 5-month moving average of each model’s accuracy on MC questions (visualized in Figure 4), we start by fitting a linear regression line on the first 10 months of data. We then add an additional month and compute a new regression on the larger window, repeating until we reach the final month, and applying an exponential decay weighting to past data to reduce the influence of distant observations. With this, we can analyze how the slope of our regression line changes as each month is added to the data. The slope in each case is negative after the cutoff data and for Claude-3.5-Sonnet, GPT-4, and Llama-3-8B, the slope eventually or immediately becomes more negative than it was at any point preceding the cutoff. Both Claude-3.5-Sonnet and Llama-3-8B have a crossover from positive to negative slope in late summer 2022, July and August, respectively, while GPT-4’s seems to occur slightly earlier, in March of 2022. For GPT-3.5, GPT-4, and Llama-3-8B, the slope becomes increasingly negative not long after the knowledge cutoff, giving evidence for a rapid decline in the near future. Likewise, the period preceding the cutoff shows a less negative slope and in some cases, most obviously in the case of GPT-3.5 and Llama-3-8B, a slightly negative but consistent slope, in other words, a gradual decline in the recent past.
(a) (a) Claude-3.5-Sonnet
(b) (b) GPT-4
(c) (c) Llama-3-8B
(d) (d) GPT-3.5
@@
Figure 7: Coefficients for regressing accuracy on the MC questions against time, as the number of months grows. Using an initial window of 10 months, we progressively add data for additional months to our regression and plot the coefficient (slope) for the regression of accuracy against time. For our regression, we use the moving average of accuracy and apply exponentially decaying weights to older months (i.e. given a window of $k$ months, we weight $x_{t}$ with $\lambda^{k-t}$; in this case $\lambda=0.995$). In each case, the slope becomes more negative following the knowledge cutoff period compared to the months immediately preceding it.
Need for Continuous Pretraining.
@@
The overall decline trend may come from two sources, the missing knowledge of future and a lack of up-to-date language representation. While the lack of knowledge can be partially recovered with information retrieval as shown in the constrained open-book setting, the gold article setting provides an “upper bound” of open-book retrieval. When provided the gold articles, the remaining decline in the model’s performance suggests continuous pre-training of LLMs (Jang et al.,, 2022; Jin et al.,, 2022; Ke et al., 2022a, ; Ke et al., 2022b, ; Yıldız et al.,, 2024) is still needed in the context of news event forecasting to address outdated representations.
@@
Interestingly, Figure 4 reveals a higher rate of performance decline around September 2021, which is the knowledge cutoff date of GPT-3.5, across all models, particularly for MC questions. In contrast, performance remains relatively stable prior to this date. We hypothesize that this trend arises because the period up to September 2021 may be overrepresented in many pretraining corpora (Raffel et al.,, 2020; Gao et al.,, 2020; Kobayashi,, 2018; Gokaslan and Cohen,, 2019; Zhu,, 2015; Rae et al.,, 2019; Tiedemann,, 2016; Saxton et al.,, 2019), compared to more recent periods. Another potential cause of this imbalance is an increasing number of websites restricting access to web crawlers (Longpre et al.,, 2024) after the rise of ChatGPT.
@@
On the data generation side, the generated questions as well as the distractor answers could contain biases from an outdated LLM, making the benchmark less reliable in the long run unless we upgrade the models. Additionally, generating questions from news articles can introduce bias by focusing only on events that have definitively occurred or not occurred. This approach overlooks speculative events, which may be of interest in forecasting questions but may not be covered in news articles, limiting the scope of the dataset. On the evaluation side, our paper proposes the continuous evaluation benchmark but at the time of the writing there isn’t a long enough time horizon on each model, especially after the cutoff dates, for a thorough analysis. Ideally, we would like to analyze the relation between the effect of knowledge and RAG cutoff dates but the trend seems to be weak within the time horizon available.

### v2

@@ 4 Experiments
We first introduce three evaluation settings in Section 4.1: 1) no access to external information, 2) access to retrieved recent news articles, and 3) access to gold articles. Section 4.2 presents the results, and Section 4.3 provides deeper insights into the observed degradation patterns.
Figure 3: Results for the closed-book setting. We plot the 5-month moving average accuracy for TF and MC questions across various models, showing LLMs’ performance degradation in future event prediction.
Table 3: For evaluating various LLMs with different knowledge cutoffs (K-Cutoffs), we show the yearly average accuracy (calculated as the average across months) from 2020 to 2024, along with the average YoY accuracy change (%) before the knowledge cutoff date (Pre-Cutoff), after the knowledge cutoff date (Post-Cutoff), and the overall average YoY accuracy change across all months (Avg).
2020 2021 2022 2023 2024 Pre-Cutoff Post-Cutoff Avg
TF Claude-3.5-Sonnet Apr 2024 81.21 79.88 78.05 74.38 64.29 -4.77 -12.41 -5.58
GPT-4 Apr 2023 69.68 66.41 60.36 60.54 56.90 -5.83 -1.96 -4.75
GPT-3.5 Sept 2021 62.86 60.12 59.36 57.11 56.09 -4.33 -3.43 -2.84
Mixtral-8x7B Unknown 57.83 52.69 43.09 39.34 36.01 -10.78
Mistral-7B Unknown 57.57 54.65 48.22 41.35 40.92 -7.75
Llama-3-8B Mar 2023 65.06 64.24 62.35 58.68 56.99 -1.95 -6.50 -2.97
Qwen-2-7B Unknown 62.42 60.18 57.67 53.39 53.04 -3.75
Gemma-2-2B Jul 2024 58.71 59.31 57.64 56.61 55.79 -1.41 -3.68 -1.04
MC Claude-3.5-Sonnet Apr 2024 76.86 77.67 74.32 69.37 61.84 -6.26 -11.78 -5.03
GPT-4 Apr 2023 70.60 70.62 66.76 56.36 51.63 -4.23 -18.54 -7.04
GPT-3.5 Sept 2021 50.27 50.40 44.38 41.45 43.09 0.14 -0.31 -3.08
Mixtral-8x7B Unknown 57.38 56.97 50.76 47.10 46.31 -4.68
Mistral-7B Unknown 50.07 52.36 48.06 44.40 42.99 -2.82
Llama-3-8B Mar 2023 52.44 54.18 50.66 47.94 46.95 -2.21 -1.25 -2.30
Qwen-2-7B Unknown 55.28 55.93 53.44 49.77 49.37 -2.35
Gemma-2-2B Jul 2024 47.87 50.71 46.81 45.20 43.28 -4.46 -4.07 -1.98
@@ 4.1 Experimental Setup
In addition to a closed-book evaluation, we explore the constrained open-book setting: how access to news articles up to different time cutoffs influences LLM performance using RAG (Lewis et al., 2020). We introduce the concept of the RAG cutoff (R-Cutoff), which limits the latest accessible date for retrieving articles. To prevent the models from leveraging information beyond the resolution date, for any question with a resolution date ($d_{\text{res}}$), the accessible articles span from January 1st, 2019 (the start of our news corpus) up to whichever comes first between the day before the resolution date and the RAG cutoff date ($d_{\text{R-Cutoff}}$). Formally, the accessible date range is $[01/01/2019,\min(d_{\text{res}}-1,d_{\text{R-Cutoff}}))$. Following prior work (Jin et al., 2021; Zou et al., 2022; Zhang et al., 2024), we employ BM25 (Robertson et al., 1995) as the retriever and select the top 5 articles relevant to each question. We truncate each retrieved article to a maximum length of 512 words. These articles are then incorporated into the input prompts to serve as additional information.
@@
We further include a setting where models are provided direct access to the gold article, from which the question is generated. [^7] See Appendix B.5 for a case example of evaluating LLMs under all three different settings. This transforms the forecasting questions into reading comprehension ones, which can also access LLMs’ general question-answering capabilities. Achieving high accuracy here ensures that the questions from our Daily Oracle dataset are answerable.
@@
Accuracy score is used as the evaluation metric. Though LLMs are tested daily, to show clearer trends, we plot the monthly performance in Figure 3, and apply a 5-month moving average to smooth the curve. We also report yearly averages and average year-over-year (YoY) accuracy change before and after models’ knowledge cutoff dates in Table 3. Additionally, despite prompting the models to avoid responses like “I cannot predict the future” and instead provide definitive answers, there are cases where such refusals still occur. The refusal rates are provided in the Appendix B.2, and these cases are counted as incorrect to ensure comparability across model results.
@@ 4.2 Main Results
Figure 3 and Table 3 present our primary results for the closed-book setting. The “Avg” column in Table 3 shows the average YoY accuracy change of all months, revealing a clear degradation in performance over time across all models on both TF and MC questions. When comparing accuracies from the beginning to the end of the evaluation period, we observe that, on average, the models’ performance declines by 21.55% on TF questions (from 64.68% to 50.74%) and by 11.33% on MC questions (from 58.30% to 51.69%). This indicates that while LLMs demonstrate certain abilities to understand real-world events and make predictions, they struggle to maintain these abilities.
Notably, the average YoY accuracy declines provide further insight. Before the knowledge cutoff, the average YoY decline across all models was relatively moderate. However, post-knowledge cutoff, we observe steeper declines in many models, with GPT-4 showing the most drastic drop in MC performance, declining by 18.54%, compared to just 4.23% before the cutoff. This contrast highlights that while LLMs manage to retain a baseline of past knowledge with small degradation, their ability to forecast future events deteriorates much more rapidly as they move beyond their training data, struggling with temporal generalization.
Among different models, Claude-3.5-Sonnet (Anthropic, 2024) significantly outperforms all others, while GPT-4 excels in MC questions but its performance in TF is not as remarkable as in MC. GPT-3.5, Qwen-2-7B (Yang et al., 2024) and Llama-3-8B (Dubey et al., 2024) show smaller temporal declines than GPT-4 in both TF and MC questions. Interestingly, Mistral-7B (Jiang et al., 2023) and Mixtral-8x7B (Jiang et al., 2024) show the most pronounced drops in TF accuracy, with scores falling below the random baseline 50% due to increased answer refusals, as shown in Figure 9. Gemma-2-2B (Team et al., 2024) exhibits the most consistent performance with the smallest average YoY decline, likely due to its more recent knowledge cutoff date.
@@
In Figure 4, we present the results of the constrained open-book setting, with Mixtral-8x7B on TF questions and Llama-3-8B on MC questions across different RAG cutoff dates. [^8] Refer to Appendix B.4 for results of other models in the constrained open-book setting. For Mixtral-8x7B, as the RAG cutoff dates extend to closer to the resolution dates, we observe a clear improvement in performance, indicating the model benefits from increasingly updated information retrieval. However, there are noticeable performance drops immediately after each RAG cutoff date when compared to providing information up to the day before the resolution date. This highlights the importance of keeping up-to-date information for optimal RAG performance. Interestingly, RAG does not uniformly enhance performance. Llama-3-8B may perform worse than the closed-book setting when the RAG cutoff is prior to the knowledge cutoff dates, suggesting outdated information may negatively impact performance. Conversely, for more recent RAG cutoff dates that extend beyond the knowledge cutoff, significant performance improvements are observed (as illustrated by the curves with cutoffs in September 2023 and March 2024). Notably, across all different RAG cutoffs, the overall performance decline pattern persists, likely due to outdated internal representations and the model’s inherent knowledge limitations.
@@
Figure 5 shows that when given access to the gold articles from which the questions are generated, LLM performance can approach around 90%, demonstrating the answerability of Daily Oracle. [^9] Results for GPT-3.5 are provided and discussed in Appendix B.3, as this older model performs relatively poorly and including it on the same scale would obscure the trends of other models. However, most of the models still show declining trends. This is noteworthy because, ideally, LLMs are expected to achieve consistent accuracy regardless of the article’s publication date when answers are directly accessible. However, the outdated representations hinder their ability to consistently generate correct answers, even in a reading comprehension setting.
@@ 4.3 Discussion
LLMs’ Performance Degradation Pattern Over Time.
@@
We observe LLMs’ performance evolution patterns in Figure 3: (1) Gradual Decline in the Recent Past: In the months before the knowledge cutoff date, which we call the recent past, we observe a gradual decline in model performance, as seen in Llama-3-8B, GPT-4, and Claude-3.5-Sonnet, likely due to a lack of representation of recent news in the training data. (2) Rapid Decline in the Near Future: In the near future, which we define as the months following a model’s knowledge cutoff date, sharp performance drops are observed in several models in MC questions. For instance, the decline in Claude-3.5-Sonnet and GPT-4 accelerates soon after their knowledge cutoffs. Most of the models, however, do not lose all the predictive power at once, as evidenced by the further decline into the farther future.
@@
We explore this further by analyzing the slope of accuracy as a function of time. In Figure 6, we show how the slope changes as we fit a regression to an increasingly larger window of data, until we reach the full set of accuracies. Specifically, using the 5-month moving average of each model’s accuracy on MC questions (visualized in Figure 3), we start by fitting a linear regression line on the first 10 months of data. We then add an additional month and compute a new regression on the larger window, repeating until we reach the final month, and applying an exponential decay weighting to past data to reduce the influence of distant observations. With this, we can analyze how the slope of our regression line changes as each month is added to the data. The slope in each case is negative after the cutoff data and for Claude-3.5-Sonnet, GPT-4, and Llama-3-8B, the slope eventually or immediately becomes more negative than it was at any point preceding the cutoff. Both Claude-3.5-Sonnet and Llama-3-8B have a crossover from positive to negative slope in late summer 2022, July and August, respectively, while GPT-4’s seems to occur slightly earlier, in March of 2022. For GPT-3.5, GPT-4, and Llama-3-8B, the slope becomes increasingly negative not long after the knowledge cutoff, giving evidence for a rapid decline in the near future. For example, Llama-3-8B’s slope is around -0.06% per month near its cutoff in March 2023, but declines by more than 2 times to approximately -0.13% per month by the end of 2023. Likewise, the period preceding the cutoff shows a milder decline, with models like GPT-3.5 and Llama-3-8B exhibiting slightly negative but consistent slopes (approximately -0.01% and -0.06% per month respectively, over the four months leading up to the cutoff). This suggests a gradual decline in the recent past.
@@
Figure 6: Coefficients for regressing accuracy on the MC questions against time, as the number of months grows. Using an initial window of 10 months, we progressively add data for additional months to our regression and plot the coefficient (slope) for the regression of accuracy against time. For our regression, we use the moving average of accuracy and apply exponentially decaying weights to older months (i.e., given a window of $k$ months, we weight $x_{t}$ with $\lambda^{k-t}$; in this case $\lambda=0.995$).
Need for Continuous pre-training.
@@
The overall decline trend may come from two sources, the missing knowledge of future and a lack of up-to-date language representation. The absence of relevant future information can lead to two outcomes: either the model makes uninformed or incorrect predictions, or, in some cases, more likely to refuse to answer altogether. We observe this latter behavior notably in Mistral-7B and Mixtral-8x7B, where refusal rates are significantly higher compared to other models, as shown in Figure 9(b). [^10] See Appendix B.2 for more discussion of the refusal behavior. The lack of knowledge can be partially recovered with information retrieval, as seen in the constrained open-book and gold article settings. For instance, Figures 9, 10, and 11 show that Mixtral-8x7B’s refusal rate drops from 14–28% in closed-book to 3–15% with open-book retrieval, and further to 0.5–4.2% with gold articles. However, accuracy still declines over time. Notably, the gold article setting provides an “upper bound” of open-book retrieval. The remaining performance drop despite full access to relevant information suggests that the models’ internal representations are outdated. This indicates continuous pre-training of LLMs (Jang et al., 2022; Jin et al., 2022; Ke et al., 2022a, b; Yıldız et al., 2024) is still needed in the context of news event forecasting.
@@
Interestingly, Figure 3 reveals a higher rate of performance decline around September 2021, which is the knowledge cutoff date of GPT-3.5, across all models, particularly for MC questions. In contrast, performance remains relatively stable prior to this date. We hypothesize that this trend arises because the period up to September 2021 may be overrepresented in many pre-training corpora (Raffel et al., 2020; Gao et al., 2020; Kobayashi, 2018; Gokaslan & Cohen, 2019; Zhu et al., 2015; Rae et al., 2020; Tiedemann, 2016; Saxton et al., 2019), compared to more recent periods. Another potential cause of this imbalance is an increasing number of websites restricting access to web crawlers after the rise of ChatGPT (Longpre et al., 2024).
@@
On the data generation side, the generated questions as well as the distractor answers could contain biases from an outdated LLM, making the benchmark less reliable in the long run unless we upgrade the models. Additionally, generating questions from news articles can introduce bias by focusing only on events that have definitively occurred, overlooking potential events that never occur and thus never appear in the news. On the evaluation side, our paper proposes the continuous evaluation benchmark but at the time of the writing there isn’t a long enough time horizon on each model, especially after the cutoff dates, for a thorough analysis. Ideally, we would like to analyze the relation between the effect of knowledge and RAG cutoff dates but the trend seems to be weak within the time horizon available.